OpenStudy (anonymous):

The amplitude and period of y = cos3xcos7x + sin3xsin7x are ?

OpenStudy (anonymous):

I was reading that lol :)

OpenStudy (anonymous):

cos3xcos7x + sin3xsin7x =cos(7x-3x)=cos4x for cos function, amplitude =1 therefore cos4x, amplitude is the same which is 1 period of cosx = 360 that of cos 4x = 360/4 = 90

OpenStudy (anonymous):

I have some amendment. Sorry for that

OpenStudy (anonymous):

why is period of cos x = 360 ?

OpenStudy (anonymous):